Can you please help me with this? Steve's Tree Farm sells Douglas firs and noble firs. One November they sold 179 more Douglas firs than noble firs. The total number of trees sold was 499. What equation could be used to solve for n, the number of noble fir trees sold? Here is my answer, but I don't think it is correct.
2n+179=499

Answer by checkley77(12844)

2n+179=499 is the correct equation.
n+(n+179)=499
2n+179=499
2n=499-179
2n=320
n=320/2
n=160 number of noble firs.
160+179=339 number of douglas firs.
proof:
160+339=499
499=499